# Alleviating Non-identifiability: a High-fidelity Calibration Objective for Financial Market Simulation with Multivariate Time Series Data

### v2

@@ 5 Conclusions
This paper studies the non-identifiability issue of social simulation. To our best knowledge, this is the first time that the non-identifiability is formally defined and analyzed to be alleviated. It is defined as the probability of randomly sampling in the non-identifiable set of the parameter space. It is rigorously analyzed that this probability can be reduced with more time series features in an exponential manner. Based on the analysis, a new objective function is proposed to effectively aggregate multiple features with a maximization function over existing discrepancy metrics. Extensive empirical studies have been conducted on 10 synthetic data and 1 real data of the financial market simulation tasks. It has been successfully verified in the simulations that, with multiple features, not only the non-identifiable set can be minimized but the simulation fidelity is largely improved. Hence, this paper provides a high-fidelity objective function to calibrate the high-frequency financial market simulator.

### v3

@@ 5 Conclusions
This paper studies the non-identifiability issue of social simulation. To our best knowledge, this is the first work that the non-identifiability is formally defined and theoretically alleviated. It is defined as the probability of randomly sampling in the non-identifiable set of the parameter space. It is rigorously analyzed that this probability can be reduced with more time series features in an exponential manner. Based on the analysis, a new objective function is proposed to effectively aggregate multiple features with a maximization function over existing discrepancy metrics. Extensive empirical studies have been conducted on 10 synthetic data and 1 real data of the FMS tasks. It has been successfully verified in the simulations that, with multiple features, not only the non-identifiable set can be minimized but the simulation fidelity is largely improved.